11) Each week a soft drink machine sells x cans of soda for $075/soda. The cost to the owner of the soda machin or each soda is So.10. The weekly fixed cost for maintaining the soda machine is $25/week. Write an equation tha relates the weekly profit, P, in dollars to the number of cans sold each week. Then use the equation to find the weekly profit when 92 cans of soda are sold in a week. A) P 0.65x-25; $34.80 C) P 0.75x-25; $44.00 B) P-0.75x+25; $94.00 D) P 0.65x+25; $84.80
Answer #1

Answer : A)

Selling price = $0.75

Cost price = $0.10

Weekly cost of maintain soda machine = $25

Number of cans sold each week = x

Profit = P = price earned from selling - cost price - fixed weekly cost

P = 0.75x - 0.10x - 25 = 0.65x-25

For 92 cans,

P = 0.65*92-25 = $34.80